OpenStudy (voltage):

Will fan and medal. Part A: If (6^2)x = 1, what is the value of x? Explain your answer. Part B: If (6^0)x = 1, what are the possible values of x? Explain your answer.

OpenStudy (photon336):

Is this what you mean \[6^{2x} = 1 \]

OpenStudy (photon336):

@Voltage

OpenStudy (voltage):

yep

OpenStudy (photon336):

we could probably use logs familiar?

OpenStudy (voltage):

ok

OpenStudy (agent0smith):

@Photon336 you don't need logs for these two

OpenStudy (voltage):

So?

OpenStudy (photon336):

\[6^{2x} = 1 | 6^{2*0} = 1 \]

OpenStudy (photon336):

so it's this \[6^{0*x} = 1 \]

OpenStudy (voltage):

Ok

OpenStudy (photon336):

, thanks for pointing that out @agent0smith actually if you take the logs of both sides you get log(1) = 0 so you don't need to use them.

OpenStudy (voltage):

So then If I multiply 6^2 times 1/6^2, ill get 6^2/6^2 which is equal to 1?

OpenStudy (voltage):

??? Hello?

OpenStudy (photon336):

So this is what you're saying \[36*(\frac{ 1 }{ 6^{2} }) = \frac{ 36 }{ 36 } = 1 \]

OpenStudy (voltage):

Yep

OpenStudy (photon336):

hmm \[6^2x = 36^{x}\]

OpenStudy (photon336):

I think there's only one answer to this \[36^{x} = 1 \] x would still be zero

OpenStudy (voltage):

Can you maybe kind of explain this im having trouble understanding

OpenStudy (photon336):

yeah no problem

OpenStudy (photon336):

I believe the question is based off this fact: \[a^{x} = 1 \] \[a^{0} =1 \]

OpenStudy (photon336):

that when we raise any number to the zero power we get 1

OpenStudy (voltage):

ok

OpenStudy (photon336):

okay I found an alternative explanation it's a little more indepth

OpenStudy (photon336):

say this is the law for dividing exponents \[\frac{ n^{x} }{ n^{y} } = n^{x-y}\]

OpenStudy (voltage):

ok

OpenStudy (photon336):

first what's \[\frac{ 5^{3} }{ 5^{2} } = ? \]

OpenStudy (voltage):

Well since they both have the same base and we are dividing we would need to subtract both of the exponents so our answer should be 5^1

OpenStudy (voltage):

I think

OpenStudy (photon336):

so now' what if we had something like this \[\frac{ 5^{3} }{ 5^{3} } = 5^{3-3} = 5^{0}\] we know that from the left hand side we've got this \[\frac{ 5^{3} }{ 5^{3} } = 1 \] now putting this together from the left and right hand side is this. \[1 = 5^{0}\]

OpenStudy (voltage):

I don't get it

OpenStudy (photon336):

look at the left hand side what is 5^3/5^3

OpenStudy (voltage):

Umm 5^0?

OpenStudy (photon336):

the argument is 1 = 5^0

OpenStudy (photon336):

\[so~if~we~take~any~number |n^{0} = 1 \]

OpenStudy (voltage):

?

OpenStudy (photon336):

\[\frac{ n^{1} }{ n^{1} } = n^{1-1} | n~is~any~real~number | 1 = n^{0}\]

OpenStudy (voltage):

ok

OpenStudy (agent0smith):

This fact should be drilled into your head. Anything to the power of 0 is 1. Eg. \[\Large \left( 55x^3 y^2 z^9 \right)^0 = 1\]

OpenStudy (voltage):

ok so

OpenStudy (agent0smith):

So... hopefully that fact gets drilled into your head. If there is a zero exponent on something, then it is equal to 1.

OpenStudy (voltage):

ok i got the answer for part A Thanks